Sol Summary – March 12th

Crew 206 March 12th 2019
Sol 16

Summary Title :

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : This morning we managed to get up a little earlier and to start the workout session on time. The planned EVA went fine. The purpose was to change the LOAC’s battery and to continue training the non-veteran crew members, two of them will be commanders next year, so it is very important to have every detail of a MDRS mission mastered. They went to white rock canyon to do some exploration and take some footage. Gaspard’s app also worked today : he is developing an app to improve EVA tracking by allowing the leader to enter data on his phone only by clicking on the screen (with the glove). This app was developed by Louis Mangin (Crew 189 Commander) last year, and this app was further developed by Gaspard this year. This afternoon we didn’t have any TELEOP session so we took some time for ourselves. More than half of the crew took a nap, and we are feeling better tonight. The VR videos we watched are quite impressive, beautifull and relaxing. The crew is feeling well this last week and will work as hard as the two other weeks.

Look ahead plan : In the next few days, for our last EVAs we will try to explore some places we haven’t explored yet, and start analysing some of the collected data we can analyse.

Anomalies in work : None

Weather : Cloudy all day

Crew physical status : Fine

EVA : LOAC Maintenance and White Rock Canyon exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Request, EVA Report

Support Requested : None

GreenHab Report – March 12th

Crew 206 Green Hab Report – 12-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (4 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.9°C

High: 32.4°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 3.2 gallons

Water in Blue Tank: 212.8 gallons

Crops watered at 0910 hours (1.1 gallons) and 1830 hours (2.1 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 08.30 AM 21.0° C / 34 %
2) 12.00 AM 29.5° C / 21 %
3) 15.00 PM 24.1° C / 24 %
4) 18.30 PM 19.7° C / 35 %

Changes to crops: Tomatoes keep sprouting. New peas were planted.

Narrative:
Today 12 peas were planted, in 4 pots.
I did my experiment in the afternoon and made interesting obervation. See Science Report for more information.

Harvest: 10 g of radish, 14 g of basil and 3 g of lemon basil.

Support/supplies needed: None.

Journalist Report – March 12th

Crew 206 – 03/12/2019

Sol 16

Authors : Benjamin Auzou, Journalist – Aurélien Murge, Astronomer and Executive Officer

"The force of the experience"

Our Association has been sending crew to the MDRS for five years. MDRS 151,164,175,189,206 are a ‘succession’ of experiences given each year to the new members. 22 French aeronautics and aerospace engineering students from ISAE-SUPAERO have worked to forge our knowledge about Martian simulations. At the beginning two veterans were part of the next crew with a shared leadership. Our crew is commanded by Jérémy, former GreenHab Officer. To continue the partnership with the MDRS and as we had this year in our association more than 30 applications for 2020 missions, the Station Director agreed to receive two French crews.
The whole crew is still undergraduate and next year will be in internship or universities all around the world. Two of us will command these crews but we can’t know yet who it will be. This is why the entire crew is trained to be commander. Everyone is taking in charge some communication windows, leading EVA’s, participating to the writing of the mission report and leading scientific experiments. Moreover, the whole crew is discussing about roles so that each member of the crew is aware of the basic knowledge of every role to be able to train the following crews.
That’s how our crews are based : experience. The members of the crews are selected, formed and advised by members of the previous missions. This allows us to have a good knowledge of the MDRS fundamentals and to deploy scientific experiments over the years. For example the aerosol counter (LOAC) analyzed the air of the desert in 2018 and in 2019 as well as the rover driving experiment ‘TELEOP’. We developed an app to help the leader of an EVA to take notes and follow protocols (created by Louis Mangin, Commander of crew 189 and improved by Gaspard, our crew engineer). And next year we will continue to work on improvement of our monitoring, but also take new experiments to the MDRS with two crews composed of talented and curious students from ISAE-SUPAERO. If you want to follow the preparation of the future missions of our association : call for projects, experiments, presentation of the crews, you can visit our Twitter account : @MDRSSupaeroCrew.

This week we continued to film moments of life and science to prepare the mission video. This video will be ready a few days after the end of the mission. What we are living for three weeks is incredible and the result of our year of preparation, and we are eager to share what we achieved with our relatives and all the people that follow our adventure. Here on Mars, our role is mainly scientific, but back on Earth our role will be more about popularizing our analog mission to children, to our friends, to our families.
These are the two goals of our association, the Club MARS from ISAE-SUPAERO : science for space and link with the public.

Operations Report – March 12th

Crew 206 Operations Report 12 – MAR – 2019

SOL: 16

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 13 hours

Solar— SOC % – Turned off at 100%, Still off around 80% (1930)

Diesel Reading – 60%
Propane Reading – 75%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 340 gallons

Water differential (static tank) – 12 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 213 gallons

Static to Loft Pump used – Yes

Water Meter: 0141745,8 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system being repared.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used

Hours: 80,5 hours

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Spirit rover used: Used

Hours: 81,9 hours

Beginning charge: 100%

Ending charge: 44%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the pipeline has a leak. Atila is working on it, more informations will follow.
Curiosity showed an engine issue light during this morning’s EVA, stopping it from functioning. Restarting the rover seems to have overcome the issue for the end of the EVA.
Curiosity has been left turned on after this morning’s EVA on mistake, but it was plugged in.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not Used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

EVA Report – March 12th

Crew 206 EVA Report 12-MAR-2019
EVA #12

Author: Gaspard Thieulin

Purpose of EVA:
After the wind of past days, the LOAC is back in the Hab for inside measurements but we need to redeploy it outside for the last week. We updated the water protection for these rainy days.

Start time: 09:25 am

End time: 11:14 am

Narrative:
This EVA was quite rich in goals, and we successfully fulfilled every one of them. First we headed up to the LOAC and the weather station, using Spirit to drive Norbert and me, and Curiosity with Benjamin and Aurelien and all our heavy stuff. We had to switch the battery of the LOAC and check on the weather station.
To perform the battery switch, I used our "EVA tracking" app, and its new protocol mode. Following simple vocal instructions, we were able to replace the battery without depending on HabCom giving us instructions. While doing so, I noticed the hygrometry sensor of the station was out of batteries, which we will change on the next EVA.
After the first step of the morning, we took back the rovers to go to White Rock Canyon at 0943. We had to use Spirit, knowing its batteries are weaker than the others, so we were very careful about our driving all the way there. In the end, we were able to stop at our planned location with 74% left on spirit and more than 80% on Curiosity at 0958.
We went for a little walk inside and around the Canyon, fulfilling the second goal of the EVA : taking footage for the after-mission movie. While doing so, we had a funny incident with Norbert who found a deep water pound by mistake : his foot went down 50cm according to his words ! Using the "EVA tracking" app, I was even able to note this happened at precisely 1028.
We ended both the walking and the fiming with rovers shots at 1053. We then headed back to the Hab, again being extremly careful to our driving to save some Spirit’s battery, especially knowing it goes down much faster after 70%. Indeed, Spirit arrived to the bar of 60% at 1102, still far from the Hab. On the road, we also had a little issue with Curiosity, stopping for what appeared to be an engine issue, at 1107. Restarting the rover allowed us to safely arrive back to the MDRS at 1110.
In the end, I’m able to say the last purpose of the EVA, proving an "EVA tracking" app can work and be useful, is a success, providing me with detailed data on our EVA.

Destination: LOAC and Weather Station location, White Rock Canyon

Coordinates: 518500/425700 (LOAC and Weather Station), 52200/4247400 (White Rock Canyon)

EVA Participants: Gaspard Thieulin (NG, leader), Benjamin Auzou (COMMS), Aurélien Mure (XO), Norbert Pouzin (GHO)

Roads and routs per MDRS Map: Entrance Road, Cow Dung Road (South)

Mode of travel: Rovers and Walking